# i-PI 3.0: a flexible, efficient framework for advanced atomistic simulations

### v1

@@ II.4 Main Features
stochastic velocity rescaling,[50] path-integral Langevin equation thermostats,[51] generalized Langevin equation (GLE) thermostats, including the optimal sampling,[52, 53] quantum, [54], $\delta$,[55] hot-spot, [56] and PIGLET thermostats,[57] dynamical corrections to Langevin trajectories,[58] fast forward Langevin thermostat,[59] Langevin sampling for noisy and/or dissipative forces.[60, 61]
@@
steepest descent, L-BFGS, trust radius BFGS, etc, nudged elastic band [66, 67, 27] and mode following saddle point search.[68]
Vibrational analysis:
@@
thermodynamic integration,[70], metadynamics, replica exchange MD,[71] and quantum alchemical transformations.[72, 73]
@@
ring polymer molecular dynamics (RPMD),[74, 75] centroid molecular dynamics (CMD),[76, 77] thermostatted RPMD,[78, 58] path integral coarse graining,[79] ring polymer instanton rates and tunnelling splittings.[80]
@@
fourth-order path integrals,[81] perturbed path integrals,[82] open path integrals,[83] the scaled-coordinate heat capacity estimator,[84] isotope fractionation estimations,[85] particle momentum distribution,[86] reweighted fourth-order path integral MD.[87, 88]
@@
Molecular dynamics driven by Electronic Structure Quantities:
non-adiabatic tunnelling rates in metallic systems,[97, 98] molecular dynamics with with time-dependent external fields.
Committee MLIP:
Figure 2: Simplified picture of the impact of a given overhead $\tau_{\text{O}}$ on the simulation throughput (expressed in ns/day, assuming a time step of 1 fs) as a function of the force evaluation cost. An overhead of about 10 ms (typical for a moderately complicated simulation with i-PI 2.6) is negligible when used with ab initio force evaluations, but becomes significant for the most efficient (or heavily-parallelized) machine-learning potentials.
Figure 3: (a) System size dependence of the timing per step when performing a NVE simulation of an ideal gas using LAMMPS, and when using i-PI with a dummy driver that returns zero energy and forces, communicating through UNIX sockets. The overhead relative to LAMMPS is also broken down into (b) a base Python overhead (obtained by repeating the i-PI simulation with no forcefield and subtracting the LAMMPS timings) and a remainder, that is associated with the communication overhead (c). Markers correspond to actual simulations while the dashed lines are fits to $\tau_{\text{O}}=A+BN_{\text{a}}$, where $N_{\text{a}}$ is the number of atoms.
@@ III Efficiency of large-scale simulations
i-PI’s socket-based communication and Python architecture offer convenience and flexibility to implement new and advanced simulation algorithms, facilitating integration with other codes. This flexibility, however, comes at the price of a computational overhead associated with UNIX- and TCP-socket communications, and Python’s inherent inefficiency compared to lower-level languages such as Fortran, C, and C++. To analyse i-PI’s overhead on the performance of the simulations, it is instructive to separate the intrinsic wall-clock time needed to evaluate energy and forces, $\tau_{\text{F}}$, and i-PI’s overhead, $\tau_{\text{O}}$, such that the length of simulation that can be obtained for a certain computational effort is proportional to $1/(\tau_{\text{F}}+\tau_{O})$. The initial focus of i-PI was on ab initio force providers, where $\tau_{\text{F}}$ is of the order of seconds, and therefore an overhead of the order of $\tau_{\text{O}}\sim 10$ ms was inconsequential (Fig. 2). However, with MLIPs becoming the primary tool for large-scale, reactive calculations, an overhead of several ms can become the performance bottleneck. Therefore, we have reevaluated i-PI’s efficiency for this release.
$A$ / ms $B$ / ($\mu$s/$N_{\text{a}}$)
@@
PIMD 2 none NPT pile_l 1.39 0.47 0.94 0.60
PIMD 2 UNIX NPT pile_l 2.62 1.27 1.32 0.76
PIMD 2 TCP/IP NPT pile_l 32.64 2.45 1.68 0.77
@@
Table 1: Comparison between the timings for the tests in the ipi_tests/profiling folder with i-PI versions 3.0-alpha2 or 2.6.3. The coefficients $A$ and $B$ can be used to estimate the overhead for a given number of atoms $N_{\text{a}}$, as $\tau_{\text{O}}=A+BN_{\text{a}}$. In the table, (PI)MD indicate simulations performed with (path integral) MD; “socket” indicates the type of communication, none corresponding to simulations that avoid entirely force evaluations. Thermostats are chosen between stocastic velocity rescaling, svr,[101] plain Langevin, and path-integral Langevin.[51] The last line corresponds to a complex setup combining path integral and replica-exchange molecular dynamics (REMD) with 4 beads and 6 ensemble replicas and a finite-difference evaluation of Suzuki-Chin (SC) high-order path integrals.[102, 81]
We decided to benchmark i-PI against the C++ based LAMMPS code, since this choice allows us to disentangle and quantify the communication and Python’s overhead and also to analyze production simulations using popular MLIPs, that are interfaced with LAMMPS. The overall cost of an i-PI simulation can be assessed by using a dummy driver that returns zero energy and forces. The communication cost can be removed by disabling the force evaluation altogether, which can be achieved by setting weight="0" as an attribute of the <force> components. Performing these benchmarks for the simplest case of constant-energy NVE MD and varying the number of atoms allowed us to determine a lower bound for $\tau_{\text{O}}$, and to subdivide it into the cost of the integrator and that of the communication (Fig. 3). It is clear that for this simple setup, with $\tau_{\text{F}}=0$, LAMMPS is almost 10 faster than i-PI 2.6 in the limit of large simulations. The relative overhead of i-PI becomes very large for a small number of atoms, where LAMMPS performs millions of steps per second, whereas i-PI 2.6 reached a hard limit of about 1 ms/step. The cost (per MD step) of these simulations can be roughly modelled as a constant term plus a per-atom contribution so that the overhead can be estimated as $\tau_{\text{O}}\approx A+BN_{\text{at}}$. Note that for such a wide range of system sizes it is not advisable to determine the coefficients by linear regression, because the intercept would be affected by large relative errors, and so we determine $A$ as the time per step of the smaller simulation (8 atoms) and $B$ as the time per step per atom of the largest one ($2^{16}$ atoms).
@@
Fig. 3 also compares the stable 2.6 version with the new 3.0 release. In the latter, we performed several optimizations to reduce the overhead associated with accessing the i-PI depend objects,[17] substituted explicit loops with array operations, and removed other obvious bottlenecks from the implementation. This reduced the overhead of the NVE integrator by a factor of 3-5, and the communication overhead by a factor of 2. Table 1 compares the value of the $A$ and $B$ coefficients for different simulation setups; in all cases, the new release dramatically reduces both constants, so that for most classical MD simulations and up to a few 1000s of atoms one can now expect an overhead of a few ms per step. It is thus possible to perform tens of millions of simulation steps per day when disregarding the force evaluation costs. The overhead is larger when considering simulations with multiple replicas, and/or complicated setups that involve e.g. high-order path integration or replica exchange. Often, however, these advanced simulation techniques allow cutting on the overall cost of a calculation by reducing the number of force evaluations to reach the desired convergence of accuracy, which is a key advantage in using i-PI over faster, but less feature-rich, alternatives. An example that is also relevant to classical MD trajectories is the presence of BAOAB integrators,[103] which allows doubling the time step in simulations of liquid water. A technical detail we also want to highlight from Table 1 is the >20×>20$\times$> 20 × reduction of the constant overhead for TCP/IP communication. This is due to the disabling of Nagle’s algorithm,[104] which was found to interfere negatively with the transfer of small messages in the i-PI communication protocol. This change has to be enacted also on the client side, so we urge developers of packages that contain an i-PI interface to apply similar changes.
@@
Figure 4: Timings for 1 step of NVT dynamics with Langevin thermostatting, for a box containing 4096 water molecules at ambient temperature and density, using a DeepMD model,[105] using the LAMMPS implementation.[106, 107] Simulations are run on nodes with 2 Intel(R) Xeon(R) Platinum 8360Y, with 36 cores each (CPU) and nodes with 2 NVIDIA V100 cards (GPU). Parallelization is based on MPI, leveraging the LAMMPS domain decomposition. (a) Timings for MD simulations. Dashed and full lines correspond to simulations performed with LAMMPS, and i-PI using LAMMPS/DeepMD as a driver, respectively. (b) Timings for PIMD simulations with 32 replicas. The shading of the curves indicates the number of independent LAMMPS tasks that are connected to i-PI, and the number of processing units indicates the sum across all tasks.
@@
Even though this release cuts significantly on the i-PI overhead, and we are committed to further reducing it in the future, it is important to determine whether and when it matters in practical use cases. To this end, we consider the task of simulating 4096 water molecules at room temperature, in the canonical NVT ensemble. We start by considering the LAMMPS implementation of a DeepMD neural network, that is used both to perform MD directly using its LAMMPS implementation,[106, 107, 7] and as a driver for both classical MD and path integral molecular dynamics with i-PI. On the hardware we use for this test (Intel(R) Xeon(R) Platinum 8360Y processors with 2 GPU NVIDIA V100 cards/node), a single-point evaluation of energies and forces on one CPU requires approximately 10s, which is several orders of magnitude longer than the corresponding i-PI overhead (approximately 2 ms/step). In practical scenarios, however, the MLIP calculation will be heavily parallelized, either using a multi-CPU implementation or using one or more GPUs. As shown in Fig. 4, the i-PI overhead remains negligible up to the highest parallelization level for CPUs (up to 4 nodes, 288 cores), even though the practical overhead at the full-node level is higher than expected because of the need of dedicating one core to the execution of i-PI. The GPU implementation of DeepMD is much faster and can be parallelized efficiently up to about 8 GPUS (4 nodes), for which (with a time step of 0.5 fs) LAMMPS clocks almost 2.5 ns of simulation per day. We observed the empirical i-PI overhead with LAMMPS DeepMD to be up to 10 times larger than that seen with the dummy driver, which we could trace to unnecessary reconstructions of the neighbor list. We then optimized the LAMMPS interface of the driver, minimizing the number of neighbor list updates and reducing the overhead to be less than 20% for the highly-efficient multi-GPU setup. For a PIMD setup, i-PI allows exploiting an additional level of parallelism by running multiple drivers in parallel. Given that this is a trivial type of parallelization, it can improve the scaling whenever the client code shows less-than-deal scaling, as is seen for the GPU implementation of DeepMD.
throughput (ns/day)
model architecture $n_{\text{PU}}$ $n_{\text{H}_{2}\text{O}}$ $\tau_{\text{F}}$/s direct w/i-PI
DeepMD CPU (Xeon) 288 4096 0.066 0.65 0.53
@@
BPNN CPU (Xeon) 768 4096 0.015 2.95 2.00
Table 2: An overview of the expected throughput for $NVT$ MD simulations of liquid water for different system sizes, hardware, and MLIPs. DeepMD simulations timings correspond to the highest level of parallelization in Fig. 4; MACE simulations use the ASE driver with the MACE-MP-0 model(S),[9] single precision and were run on one A100 GPU card. We note, however, that the ASE driver is not the most efficient way of running MACE, but regardless, the overhead of i-PI is negligible compared to a MACE force evaluation. Behler-Parrinello neural network simulations use the model from Ref. [108], and were run using the n2p2 interface in LAMMPS, on a Intel® Xeon® E5-2690 v3 @ 2.60GHz.
A simple estimate of the throughput of an i-PI calculation.
These experiments show that, in practical use cases, the overhead introduced by i-PI is negligible or small up to high levels of parallelization of the MLIP code – although one should reassess these conclusions depending on the hardware, the specific MLIP implementation, and the type of simulation performed. As shown in Fig. 2, a practical recipe to determine what throughput one can expect out of the combination of i-PI with your favourite machine-learning code is to determine the overhead $\tau_{\text{O}}$, either from Table 1 or by running a short simulation with a dummy driver. A dummy calculation is also suitable to tune the simulation parameters (and potentially to help remove bottlenecks from other simulation pathways). The cost of evaluating energy and forces $\tau_{\text{F}}$, also depends on the type of MLIP, system size and the hardware of choice. Thus, by comparing the $\tau_{\text{O}}$ and $\tau_{\text{F}}$ it is possible to determine whether the i-PI overhead is significant.
As shown in Table 2, for both small and medium-sized supercells, and for DeepMD, n2p2 and MACE models, a standard constant-temperature MD simulation is limited by $\tau_{\text{F}}$. The only exception is DeePMD(GPU), which as mentioned earlier, appears to be specific to this setup and due to reasons we continue to investigate. In summary, using i-PI as a driver not only does not prevent reaching a few million energy evaluations per day, corresponding to several ns for typical MD time steps, but also allows access to advanced simulation methods, such as path integral molecular dynamics, as well as facile interfacing with new MLIPs implementations.

### v2

@@ II.3 Main Features
stochastic velocity rescaling,[67] path-integral Langevin equation thermostats,[68] generalized Langevin equation (GLE) thermostats, including the optimal sampling,[69, 70] quantum, [71], $\delta$,[72] hot-spot, [73] and PIGLET thermostats,[74] GLE thermostats for use with path-integral correlation functions,[75] fast forward Langevin thermostat,[76] Langevin sampling for noisy and/or dissipative forces.[77, 78]
@@
steepest descent, conjugate gradient, L-BFGS, trust radius BFGS, nudged elastic band [83, 84, 34] and mode following saddle point search.[85]
@@
thermodynamic integration,[86] metadynamics, replica exchange MD,[87] and quantum alchemical transformations.[88, 89]
@@
ring polymer molecular dynamics (RPMD),[90, 91] centroid molecular dynamics (CMD),[92, 93] thermostatted RPMD,[94, 75], ring polymer instanton rates and tunnelling splittings.[95]
@@
fourth-order path integrals,[96] perturbed path integrals,[97] the scaled-coordinate heat capacity estimator,[98] isotope fractionation estimations,[99] particle momentum distribution,[100] reweighted fourth-order path integral MD.[101, 102]
II.3.2 Features added after the 2.0 release:
Efficient integrators:
BAOAB splitting [47], and Cayley integrator [103] for larger PIMD timesteps.
Barostats:
MTTK fully flexible barostat [104], MTS implementation, and barostat implementation for the Suzuki-Chin [47] PIMD.
Advanced path integral simulations and estimators:
Suzuki-Chin double virial operator estimator for calculating heat capacity  [47] and Open chain path integrals [105] for estimating momentum dependent operators
Vibrational analysis:
Particle exchange Monte Carlo:
for efficient sampling of poorly ergodic mixtures  [107]
@@
Nuclear propagation by additional dynamical variables:
non-adiabatic tunnelling rates in metallic systems,[113, 114] molecular dynamics with time-dependent external fields.
Ensemble models:
Extracting dynamical information from thermostatted simulations:
for vibrational spectra and diffusion coefficients from PIGLET thermostatted simulations [116]
@@
[…]
for approximate quantum dynamics using molecular dynamics on an MLIP representing the centroid potential of mean force [120, 121].
Figure 2: Simplified picture of the impact of a given overhead $\tau_{\text{O}}$ on the simulation throughput (expressed in ns/day, assuming a time step of 1 fs) as a function of the force evaluation cost. The larger the force evaluation time, the lower the impact a given $\tau_{\text{O}}$ on the simulation throughput. An overhead of about 10 ms (typical for a moderately complicated simulation with i-PI 2.6) is negligible when used with ab initio force evaluations, but becomes significant for the most efficient (or heavily parallelised) MLIPs.
Figure 3: (a) System size dependence of the timing per step when performing a NVE simulation of an ideal gas using LAMMPS, and when using i-PI with a dummy driver that returns zero energy and forces, communicating through UNIX sockets. The cost of a simulations run with i-PI is broken down into (b) a base Python overhead (obtained by performing the i-PI simulation with no forcefield) and the communication overhead (c), that is obtained by subtracting the Python overhead from the cost of a simulation that does communicate with the dummy driver. Markers correspond to actual simulations while the dashed lines are fits to $\tau_{\text{O}}=A+BN_{\text{a}}$, where $N_{\text{a}}$ is the number of atoms. All simulations were run on nodes with Intel Xeon E5-2690 v3 @ 2.60GHz processors using a single core.
@@ III Efficiency of large-scale simulations
i-PI’s socket-based communication and Python architecture offer convenience and flexibility to implement new and advanced simulation algorithms, facilitating integration with other codes. This flexibility, however, comes at the price of a computational overhead associated with UNIX- and TCP-socket communications, and Python’s inherent inefficiency compared to lower-level languages such as Fortran, C, and C++.
To analyse i-PI’s overhead on the performance of the simulations, it is instructive to separate the intrinsic wall-clock time needed to evaluate energy and forces, $\tau_{\text{F}}$, and the time overhead associated only to i-PI, $\tau_{\text{O}}$, such that the length of simulation that can be obtained for a certain computational effort is proportional to $1/(\tau_{\text{F}}+\tau_{O})$. The initial focus of i-PI was on ab initio force providers, where $\tau_{\text{F}}$ is of the order of several seconds, and therefore an overhead of $\tau_{\text{O}}\sim 10$ ms was inconsequential ( see Fig. 2). However, with MLIPs becoming the primary tool for large-scale, reactive calculations, an overhead of several ms can become the performance bottleneck. Therefore, we have improved the efficiency of i-PI for this release.
$A$ (ms) $B$ ($\mu$s/$N_{\text{a}}$)
@@
TCP/IP socket communication overhead 21.07 1.20 0.56 0.08
PIMD 2 none NPT pile_l 1.39 0.47 0.94 0.60
PIMD 2 UNIX NPT pile_l 2.62 1.27 1.32 0.76
UNIX socket communication overhead 1.23 0.80 0.38 0.16
PIMD 2 TCP/IP NPT pile_l 32.64 2.45 1.68 0.77
TCP/IP socket communication overhead 31.25 1.98 0.74 0.17
Table 1: Comparison between the timings for the tests in the ipi_tests/profiling directory with i-PI version 3.0.0-beta or 2.6.3. The coefficients $A$ and $B$ can be used to estimate the overhead for a given number of atoms $N_{\text{a}}$, as $\tau_{\text{O}}=A+BN_{\text{a}}$. In the table, (PI)MD indicate simulations performed with (path integral) MD; “socket” indicates the type of communication, none corresponding to simulations that avoid entirely force evaluations. “pile_l” stands for the path-integral Langevin thermostat.[68]. The purpose of this table is to make the communication overhead of MD and PIMD runs evident. The numbers are obtained by subtracting the UNIX or TCP/IP socket values from the respective line where socket is none. All the simulations were run on nodes with 2 Intel(R) Xeon(R) Platinum 8360Y processors using a single core.
$A$ (ms) $B$ ($\mu$s/$N_{\text{a}}$)
@@
Table 2: Comparison between the timings for additional tests in the ipi_tests/profiling directory with i-PI version 3.0.0-beta or 2.6.3, presented in the same way as in Table 1. In the table, (PI)MD indicate simulations performed with (path integral) MD; Thermostats are chosen between stochastic velocity rescaling, svr and [122] plain Langevin. The last line corresponds to a complicated setup combining path integral and replica-exchange molecular dynamics (REMD) with 4 beads and 6 ensemble replicas and a finite-difference evaluation of Suzuki-Chin (SC) high-order path integrals.[123, 96] All the simulations were run avoiding force evaluations entirely (socket = none), on nodes with 2 Intel(R) Xeon(R) Platinum 8360Y processors using a single core. The communication timings reported in Table 1 should be added to the timings reported here for a realistic estimation, keeping into account, when relevant, the number of replicas.
III.1 Estimating the overhead of i-PI
The cost per MD step of i-PI simulations can be roughly modelled as a constant term plus a per-atom contribution. This overhead can be estimated as $\tau_{\text{O}}\approx A+BN_{\text{a}}$. Note that, for such a wide range of system sizes, it is not advisable to determine the coefficients by linear regression because the intercept would be affected by large relative errors. Therefore, we determined $A$ as the time per step of the smaller simulation (8 atoms) and $B$ as the time per step per atom of the largest one ($2^{16}$ atoms).
We start by analysing the simplest possible case of a constant-energy simulation of an ideal gas, with no forces acting between atoms and therefore $\tau_{\text{F}}\approx 0$. (Fig. 3). In this case, the advantage of a C++ implementation such as that in LAMMPS is evident. For the largest system sizes, LAMMPS is about 10 times faster than i-PI, but for a few atoms the difference is much larger: LAMMPS can perform millions of steps per second, whereas i-PI 2.6 saturates at about 1000 steps per second (Fig. 3a).
It is also instructive to partition $\tau_{\text{O}}$ into a contribution associated with the Python overhead and the one associated with the socket communication. We estimated the former by disabling force evaluation altogether, which can be achieved by setting weight="0" as an attribute of the <force> components. We estimated the latter by subtracting the Python overhead from the overall cost of a simulation communicating with a dummy driver that simply returned zero energy and forces (Fig. 3b-c).
Fig. 3 also compares the computational efficiency of the stable 2.6 version with that of the current 3.0 release. In the latter, we performed several optimizations to reduce the overhead associated with accessing the i-PI depend objects,[18] substituted explicit loops with array operations, and removed other obvious bottlenecks from the implementation. This reduced the overhead of the NVE integrator by a factor of 3-5, and the communication overhead by a factor of 2.
In Table 1, we report the values of the $A$ and $B$ coefficients for different simulation setups enabling the calculation of the communication overhead, while in Table 2, we report the values of the $A$ and $B$ associated with the Python overhead for different nuclear motion algorithms. Thus, the total overall i-PI overhead, $\tau_{\text{O}}$, can be obtained by adding the TCP/IP or UNIX socket timings from Table 1 to the algorithm-specific times from Table 2. In all cases, the new release dramatically reduces both constants so that for most classical MD simulations and up to a few 1000s of atoms, one can now expect an overhead of a few ms per step. It is thus possible to perform tens of millions of simulation steps per day when disregarding the force evaluation costs.
@@ III.1 Estimating the overhead of i-PI
The Python overhead is larger when considering simulations with multiple replicas and/or complicated setups that involve, e.g. high-order path integration or replica exchange, as evidenced in Table 2. Often, however, these advanced simulation techniques decrease the overall cost of a calculation by reducing the number of force evaluations necessary to reach the desired statistical accuracy. This is a key advantage of using i-PI over faster, but less feature-rich, alternatives. An example, also relevant to classical MD trajectories, is the use of BAOAB integrators,[124] which allow doubling the time step in simulations of liquid water. A technical detail we also want to highlight from Table 1 is the >20×>20$\times$> 20 × reduction of the constant overhead for TCP/IP communication. This was achieved by disabling Nagle’s algorithm,[125] which was found to interfere negatively with the transfer of small messages in the i-PI communication protocol. This change has to be enacted also on the client side, as currently done in the driver codes we distribute with i-PI, so we urge developers of packages that contain an i-PI interface to apply similar changes.
@@
Figure 4: Timings for 1 step of NVT dynamics with Langevin thermostatting, for a box containing 4096 water molecules at ambient temperature and density, using a DeePMD model,[126] using the LAMMPS implementation.[127, 128] Simulations were run on nodes with 2 Intel(R) Xeon(R) Platinum 8360Y, with 36 cores each (CPU) and nodes with 2 NVIDIA V100 cards (GPU). Parallelization is based on MPI, leveraging the LAMMPS domain decomposition. Dotted lines are guides for the eye corresponding to an ideal $1/n$ scaling. (a) Timings for MD simulations. Dashed and full lines correspond to simulations performed with LAMMPS, and i-PI using LAMMPS/DeePMD as a driver, respectively. (b) Timings for PIMD simulations with 32 replicas. The shading of the curves indicates the number of independent LAMMPS tasks that are connected to i-PI, and the number of processing units indicates the sum across all tasks.
III.2 Efficient performance of i-PI on realistic parallelised simulations
Even though this release significantly cuts the i-PI overhead $\tau_{O}$, and we are committed to further reducing it in the future, it is important to determine whether and when it matters in practical use cases. To this end, we consider the task of simulating 4096 water molecules at room temperature, in the canonical NVT ensemble, using a stochastic velocity rescaling thermostat. Based on Table 1, we expect that using i-PI 3.0 should entail an overhead of 1.3 ms when using a UNIX socket, and 2.5 ms when using a TCP/IP socket, as necessary for simulations running across many nodes.
We start by considering the LAMMPS implementation of a DeePMD neural network, that is used both to perform MD directly using its LAMMPS implementation,[127, 128, 7] and as a driver for running classical MD and path integral molecular dynamics with i-PI. On the hardware we use for this test, detailed in the caption of Fig. 4, the evaluation of energies and forces on one CPU requires approximately 10 s/step, which is several orders of magnitude larger than the estimated $\tau_{O}$.
@@ III.2 Efficient performance of i-PI on realistic parallelised simulations
In practical scenarios, however, the MLIP calculation will be heavily parallelized, either using a multi-CPU implementation or using one or more GPUs. As shown in Fig. 4, i-PI and LAMMPS having essentially equivalent performance up to the highest parallelization level for CPUs (up to 4 nodes, 288 cores), consistent with the small predicted $\tau_{O}$. The GPU implementation of DeePMD is much faster and can be parallelized efficiently up to about 8 GPUS (4 nodes), for which (with a time step of 0.5 fs) LAMMPS clocks almost 2.5 ns of simulation per day. In our initial benchmarks we observed the empirical $\tau_{O}$ with LAMMPS DeePMD to be up to 10 times larger than when using the dummy driver, which we could trace to unnecessary reconstructions of the neighbor list at the LAMMPS side of the interface. We could optimize the LAMMPS interface of the driver, minimizing the number of neighbor list updates, and reducing the observed overhead to about 3 ms for the 8-GPU run. This is still almost 2 times larger than predicted, because of the additional bookkeeping that is needed to interface i-PI with a feature-rich code such as LAMMPS, in comparison to the minimalistic dummy driver. Nevertheless, these optimizations have reduced $\tau_{O}$ to less than 20% of the overall simulation step even for the highly efficient multi-GPU setup, allowing to perform more than 2 ns of simulation per day with a relatively large number of atoms. For a PIMD setup, i-PI allows exploiting an additional level of parallelism by running multiple clients in parallel. Given that this is a trivial type of parallelization, it can improve the scaling whenever the client code shows less-than-ideal scaling, as observed for the GPU implementation of DeePMD in Fig. 4b, even though the speedup is minimal in this case.
throughput
(ns/day)
model architecture $n_{\text{CPU}}$/$n_{\text{GPU}}$ $n_{\text{H}_{2}\text{O}}$ $\tau_{\text{F}}$/s direct w/i-PI
DeePMD CPU (Xeon) 288 4096 0.066 0.65 0.56
@@
BPNN CPU (Xeon) 768 4096 0.015 2.79 2.10
Table 3: An overview of the expected throughput for $NVT$ MD simulations of liquid water (with a time step of 0.5 fs) for different system sizes, hardware, number of processing units ($n_{\text{CPU}}$/$n_{\text{GPU}}$) and MLIPs. DeePMD simulation timings correspond to the highest level of parallelization in Fig. 4. The Simulations were run on nodes with 2 Intel(R) Xeon(R) Platinum 8360Y, with 36 cores each (CPU) and nodes with 2 NVIDIA V100 cards (GPU); MACE simulations use the ASE driver with the MACE-MP-0 model(S),[9] single precision and were run on one A100 GPU card. Behler-Parrinello neural network (BPNN) simulations use the model from Ref. [129], and were run using the n2p2 interface in LAMMPS, on a Intel Xeon E5-2690 v3 @ 2.60GHz.
III.3 Near-ideal performance of i-PI with popular MLIPs
These experiments show that, in practical use cases, the overhead introduced by i-PI is negligible or small up to high levels of parallelization of the MLIP code – although one should reassess these conclusions depending on the hardware, the specific MLIP implementation, and the type of simulation performed. As shown in Fig. 2, a practical recipe to determine what throughput one can expect out of the combination of i-PI with your favourite machine learning code is to determine the overhead $\tau_{\text{O}}$, either from Table 1 or by running a short simulation with a dummy driver. A dummy calculation is also suitable to tune the simulation parameters (and potentially to help remove bottlenecks from other simulation pathways– some examples of profiling runs can be found in the test directory). The cost of evaluating energy and forces, $\tau_{\text{F}}$, also depends on the type of MLIP, system size and the hardware of choice. Thus, by comparing $\tau_{\text{O}}$ and $\tau_{\text{F}}$, it is possible to determine whether $\tau_{O}$ is significant.
As shown in Table 3, for both small and medium-sized supercells, and for DeePMD, n2p2 and MACE models, a standard constant-temperature MD simulation is limited by $\tau_{\text{F}}$. Even though, as shown by the case of LAMMPS, the overhead of a specific client might be larger than that of the minimalistic driver distributed with i-PI, this type of analysis helps determine whether i-PI is a suitable choice for the modeling task, or whether one should look for (or implement) a more performant, compiled-language implementation. In summary, using i-PI as a driver does not prevent reaching a few million energy evaluations per day, corresponding to several ns for typical MD time steps. In addition, it allows access to advanced simulation methods, such as path integral molecular dynamics, as well as facile interfacing with new MLIPs implementations, which might make i-PI the preferred choice even in cases in which it entails a small performance penalty.